A test of HyperTeX

The HyperTeX extensions to TeX were designed to allow easy inclusion of html style networked hypertext within TeX. The implementation is a very straightforward application of the TeX special macros, copying the html strings directly into the dvi files, allowing an intelligent dvi previewer to understand and interpret the html-style hypertext commands. Other hypertext styles could be similarly implemented.

Definitions:

Hypertex is the (very logical) name we have given this enhanced TeX/dvi combination. We hope it will become

very popular! html is short for the Hypertext Markup Language, in principle an SGML-based hypertext system (ie. using

angle-bracket tags), which is the ”Lingua Franca” of the world-wide web. We hope HyperTeX will become an alternative for networked hypertext, particularly the kind that works better in TeX (ie. things with equations and tables and other formatted stuff). If you want to go to second page click here. Or if you want to go to second page you can also click here. Here is an external reference to a file. (if you are testing this on a machine other than dmitri.stanford.edu

you may need to change the filename in ht test.tex)

And this is another external reference. It points to the e-print database. This is a reference to a local file in current working directory. This is a reference to a local file with a given path. This is a reference to a local file with a given path starting with ///.

Second page! 2